# DualNILM: Energy Injection Identification enabled Disaggregation with Deep Multi-Task Learning

### v2

@@ A-D Signal Space Interpretation and Rank Constraints
In summary, allowing negative flows expands the solution space beyond the convex cone imposed by nonnegative appliances. This expansion degrades the ability to achieve stable solution identification and significantly increases the computational complexity of load disaggregation. Traditional NILM methods that treat all loads as nonnegative typically fail to disentangle consumption from behind-the-meter generation when measurements dip below zero or remain subdued for extended intervals, as demonstrated in our experimental results (Figure 1 and Table III).
@@ Appendix C Datasets and PV Simulation Details
C-A Simulated Datasets
C-A1 Data Alignment and Resampling
@@
C-A2 The Choice of Simulated 2kW PV Scale
For the context of REDD and UKDALE, in the early-2010s PV market, residential new-construction PV systems most commonly fell in the 2-4 kW range [46], which puts our choice of 2 kW injection at the small-system end but still representative for that period.
From a siting perspective, NREL’s 2010 residential benchmark assumes 35 m2 of array area using 27 modules (14.5% efficiency) for a 5 kWdc system; by simple proportionality, a 2 kW array needs 14 m2, well within the envelope of typical roof planes [47].
Consistently, NREL’s rooftop technical-potential study requires just one contiguous roof plane of 10 m2, noting this supports 1.5 kW at 15% efficiency, again indicating that 2 kW is feasible on modest roof areas [48].
Policy design at the time also discouraged oversizing: California’s NSHP required systems to be “sized so that the electricity produced offsets part or all of on-site needs,” with a minimum of 1 kW AC and systems 7.5 kW AC deemed “sized to serve on-site load.” This aligns with our decision not to exceed historical maximum household loads in these datasets [49].
C-A3 Appliance State Labeling Thresholds
The fixed thresholds for appliance state detection are based on typical operating powers observed in the datasets and are consistent with prior research [10, 9]. Table V summarizes the thresholds used.
@@ C-A3 Appliance State Labeling Thresholds
C-A4 Data Splitting
The datasets are split into training and testing sets based on the dates specified in Section V-A.
@@ C-A4 Data Splitting
REDD [9]: The dataset contains power sequence data for 6 US houses of 24 to 36 days between Apr. 2011 and Jun. 2011, with 1 Hz sampling frequency for the mains meter and 3 Hz for the 10 25 types of appliance meters. We utilize Houses 1–3, with different monitoring periods for each house:
@@
UK-DALE [33]: The dataset contains aggregate power consumption and measurements of 454 appliances from 5 UK houses. For house 1, the mains readings were recorded every 1 second, and appliances power readings were recorded every 6 seconds from Nov. 2012 to Jan. 2015, while for the other 4 houses, the recording periods are about half a year in 2013 or 2014. We analyze Houses 1–2, with consistent monitoring of five major appliances (kettle, microwave, refrigerator, dishwasher, and washing machine):
@@ C-A5 Simulation Details
| Washing Machine | 500.00 | 28.00 | 2.04 | 0.49 | -111There is no washing machine in REDD house2 in reality. |
@@ Appendix D Details of Experiment
D-A Details of Benchmark Methods
@@
FHMM (Factorial Hidden Markov Model): A classical probabilistic model that represents the aggregate power consumption as the sum of multiple hidden Markov chains, each modeling an individual appliance’s state [7]. We implemented FHMM using the hmmlearn library, modeling each appliance with a two-state (ON/OFF) HMM. The model was trained using the Expectation-Maximization algorithm over sequences of length 300.
XGBoost: An efficient implementation of gradient boosting decision trees capable of capturing complex nonlinear relationships in the data [41]. We utilized the XGBoost classifier with parameters set as use_label_encoder=False and eval_metric=’logloss’. The model was trained on flattened input sequences, leveraging both active power and reactive power features.
Seq2Point: A neural network architecture that maps a sequence of aggregate power readings to a point estimate of an individual appliance’s state [11]. Our implementation used a convolutional neural network with five convolutional layers:
@@
CNN-LSTM: A hybrid deep learning model combining Convolutional Neural Networks (CNNs) for feature extraction and Long Short-Term Memory (LSTM) networks for temporal sequence modeling [42]. The CNN component consisted of:
@@
Transformer: A deep learning model based on the self-attention mechanism, allowing for efficient sequence modeling without relying on recurrence [27]. We implemented a Transformer encoder with the following parameters:
@@
UNetNILM: An adaptation of the U-Net architecture for NILM tasks, designed to capture multiscale temporal patterns in the data [44]. The model featured an encoder-decoder structure with skip connections:
@@
Seq2Seq: A sequence-to-sequence neural network model that learns to map input sequences of aggregate power to output sequences of disaggregated signals [13]. Our implementation used an encoder-decoder LSTM architecture with:
@@
DAE (Denoising Autoencoder): An autoencoder model trained to reconstruct the target signal from a corrupted version of the input, effectively isolating the micro-inverter’s contribution from the aggregate power [43]. The model consisted of:
@@ Appendix E Full Experiment Results
E-A Overview of All Results Cross All Datasets
@@
Table VII and VIII show the overall average performance metrics on appliance state recognition and energy injection disaggregation tasks.
Accuracy (%) | Recall (%) | Precision (%) | F1 (%) | |||
| Laboratory Dataset (Real BTM Injection) | FHMM | 60.95 | 28.32 | 16.84 | 17.25 | |||
| XGBoost | 85.82 | 48.20 | 52.10 | 47.75 | ||||
| Seq2Point | 67.63 | 53.06 | 27.86 | 34.27 | ||||
| CNN-LSTM | 74.03 | 52.97 | 30.29 | 36.67 | ||||
| Transformer | 81.79 | 52.91 | 38.16 | 42.76 | ||||
| UNetNILM | 89.20 | 54.01 | 60.81 | 53.89 | ||||
| DualNILM | 98.54 | 85.43 | 84.16 | 84.54 | ||||
| REDD (Synthetic PV Injection) | House 1 | FHMM | 79.52 | 1.46 | 13.47 | 2.07 | ||
| XGBoost | 97.90 | 74.50 | 88.74 | 79.18 | ||||
| Seq2Point | 97.65 | 42.20 | 56.42 | 46.07 | ||||
| CNN-LSTM | 98.67 | 46.32 | 50.75 | 47.96 | ||||
| Transformer | 95.99 | 56.08 | 54.10 | 53.42 | ||||
| UNetNILM | 99.05 | 76.11 | 85.22 | 75.27 | ||||
| DualNILM | 99.84 | 96.26 | 97.11 | 96.64 | ||||
| House 2 | FHMM | 75.08 | 0.58 | 0.34 | 0.43 | |||
| XGBoost | 98.43 | 92.47 | 95.58 | 93.53 | ||||
| Seq2Point | 91.43 | 52.90 | 58.44 | 50.85 | ||||
| CNN-LSTM | 98.86 | 49.19 | 48.74 | 48.96 | ||||
| Transformer | 97.92 | 47.78 | 47.98 | 47.87 | ||||
| UNetNILM | 99.57 | 64.16 | 77.22 | 67.28 | ||||
| DualNILM | 99.70 | 92.48 | 94.00 | 93.11 | ||||
| House 3 | FHMM | 83.84 | 4.53 | 0.90 | 1.49 | |||
| XGBoost | 94.89 | 46.22 | 56.38 | 50.51 | ||||
| Seq2Point | 98.10 | 42.78 | 50.90 | 44.47 | ||||
| CNN-LSTM | 95.92 | 28.85 | 38.17 | 32.62 | ||||
| Transformer | 97.81 | 40.21 | 38.89 | 39.43 | ||||
| UNetNILM | 99.07 | 59.91 | 54.73 | 57.10 | ||||
| DualNILM | 99.69 | 77.42 | 84.46 | 80.04 | ||||
| UKDALE (Synthetic PV Injection) | House 1 | FHMM | 89.25 | 1.58 | 0.76 | 1.01 | ||
| XGBoost | 93.64 | 41.74 | 68.04 | 45.97 | ||||
| Seq2Point | 88.70 | 41.47 | 43.29 | 36.79 | ||||
| CNN-LSTM | 98.33 | 31.76 | 31.29 | 31.52 | ||||
| Transformer | 96.90 | 33.65 | 27.76 | 30.26 | ||||
| UNetNILM | 98.76 | 52.03 | 72.01 | 55.20 | ||||
| DualNILM | 99.72 | 95.23 | 94.30 | 94.71 | ||||
| House 2 | FHMM | 82.95 | 0.62 | 0.52 | 0.55 | |||
| XGBoost | 96.13 | 66.61 | 79.19 | 70.67 | ||||
| Seq2Point | 98.83 | 42.68 | 46.49 | 41.92 | ||||
| CNN-LSTM | 97.81 | 41.65 | 45.48 | 43.33 | ||||
| Transformer | 99.13 | 59.88 | 77.44 | 61.73 | ||||
| UNetNILM | 99.17 | 64.25 | 68.74 | 65.43 | ||||
| DualNILM | 99.73 | 90.63 | 94.46 | 92.31 | ||||
| Laboratory Dataset (Real BTM Injection) | FHMM | 0.5015 | 0.4492 | |||||
| Seq2Seq | 0.3512 | 0.2622 | ||||||
| DAE | 0.8830 | 0.7851 | ||||||
| UNetNILM | 0.2305 | 0.1883 | ||||||
| DualNILM | 0.1429 | 0.0804 | ||||||
| REDD (Synthetic PV Injection) | House 1 | FHMM | 0.8093 | 0.2200 | ||||
| Seq2Seq | 0.6834 | 0.2370 | ||||||
| DAE | 0.8115 | 0.2171 | ||||||
| UNetNILM | 0.7589 | 0.1983 | ||||||
| DualNILM | 0.6661 | 0.1777 | ||||||
| House 2 | FHMM | 0.8403 | 0.3367 | |||||
| Seq2Seq | 0.7217 | 0.3287 | ||||||
| DAE | 0.8448 | 0.3356 | ||||||
| UNetNILM | 2.2077 | 0.6601 | ||||||
| DualNILM | 0.7065 | 0.2768 | ||||||
| House 3 | FHMM | 0.5575 | 0.1766 | |||||
| Seq2Seq | 0.4974 | 0.1940 | ||||||
| DAE | 0.5606 | 0.1743 | ||||||
| UNetNILM | 2.2517 | 0.4708 | ||||||
| DualNILM | 0.4261 | 0.1422 | ||||||
| UKDALE (Synthetic PV Injection) | House 1 | FHMM | 1.1905 | 0.3331 | ||||
| Seq2Seq | 0.9055 | 0.2758 | ||||||
| DAE | 0.9197 | 0.2948 | ||||||
| UNetNILM | 7.5014 | 1.3871 | ||||||
| DualNILM | 0.9008 | 0.2731 | ||||||
| House 2 | FHMM | 1.1477 | 0.3521 | |||||
| Seq2Seq | 0.9252 | 0.2833 | ||||||
| DAE | 0.9594 | 0.3368 | ||||||
| UNetNILM | 1.1285 | 0.3043 | ||||||
| DualNILM | 0.9344 | 0.2883 | ||||||
E-B Laboratory Dataset Results
Table IX shows the details of the laboratory dataset performance metrics for DualNILM and benchmarks.

### v3

@@ A.4. Signal Space Interpretation and Rank Constraints
In summary, allowing negative flows expands the solution space beyond the convex cone imposed by nonnegative appliances. This expansion degrades the ability to achieve stable solution identification and significantly increases the computational complexity of load disaggregation. Traditional NILM methods that treat all loads as nonnegative typically fail to disentangle consumption from behind-the-meter generation when measurements dip below zero or remain subdued for extended intervals, as demonstrated in our experimental results (Figure 1 and Table 3).
@@ Appendix C Datasets and PV Simulation Details
C.1. Simulated Datasets
C.1.1. Data Alignment and Resampling
@@
C.1.2. The Choice of Simulated 2kW PV Scale
For the context of REDD and UKDALE, in the early-2010s PV market, residential new-construction PV systems most commonly fell in the 2-4 kW range (Barbose et al., 2014), which puts our choice of 2 kW injection at the small-system end but still representative for that period.
From a siting perspective, NREL’s 2010 residential benchmark assumes 35 m2 of array area using 27 modules (14.5% efficiency) for a 5 kWdc system; by simple proportionality, a 2 kW array needs 14 m2, well within the envelope of typical roof planes (Goodrich et al., 2012).
Consistently, NREL’s rooftop technical-potential study requires just one contiguous roof plane of 10 m2, noting this supports 1.5 kW at 15% efficiency, again indicating that 2 kW is feasible on modest roof areas (Gagnon et al., 2016).
Policy design at the time also discouraged oversizing: California’s NSHP required systems to be “sized so that the electricity produced offsets part or all of on-site needs,” with a minimum of 1 kW AC and systems 7.5 kW AC deemed “sized to serve on-site load.” This aligns with our decision not to exceed historical maximum household loads in these datasets (California Energy Commission, 2013).
C.1.3. Appliance State Labeling Thresholds
The fixed thresholds for appliance state detection are based on typical operating powers observed in the datasets and are consistent with prior research (Kelly and Knottenbelt, 2015a; Kolter and Johnson, 2011). Table 4 summarizes the thresholds used.
@@ C.1.3. Appliance State Labeling Thresholds
C.1.4. Data Splitting
The datasets are split into training and testing sets based on the dates specified in Section 5.1.
@@ C.1.4. Data Splitting
REDD (Kolter and Johnson, 2011): The dataset contains power sequence data for 6 US houses of 24 to 36 days between Apr. 2011 and Jun. 2011, with 1 Hz sampling frequency for the mains meter and 3 Hz for the 10 25 types of appliance meters. We utilize Houses 1–3, with different monitoring periods for each house:
@@
UK-DALE (Kelly and Knottenbelt, 2015b): The dataset contains aggregate power consumption and measurements of 454 appliances from 5 UK houses. For house 1, the mains readings were recorded every 1 second, and appliances power readings were recorded every 6 seconds from Nov. 2012 to Jan. 2015, while for the other 4 houses, the recording periods are about half a year in 2013 or 2014. We analyze Houses 1–2, with consistent monitoring of five major appliances (kettle, microwave, refrigerator, dishwasher, and washing machine):
@@ C.1.5. Simulation Details
| Washing Machine | 500.00 | 28.00 | 2.04 | 0.49 | -222There is no washing machine in REDD house2 in reality. |
@@ Appendix D Details of Experiment
D.1. Details of Benchmark Methods
@@
FHMM (Factorial Hidden Markov Model): A classical probabilistic model that represents the aggregate power consumption as the sum of multiple hidden Markov chains, each modeling an individual appliance’s state (Kim et al., 2011). We implemented FHMM using the hmmlearn library, modeling each appliance with a two-state (ON/OFF) HMM. The model was trained using the Expectation-Maximization algorithm over sequences of length 300.
XGBoost: An efficient implementation of gradient boosting decision trees capable of capturing complex nonlinear relationships in the data (Chen and Guestrin, 2016). We utilized the XGBoost classifier with parameters set as use_label_encoder=False and eval_metric=’logloss’. The model was trained on flattened input sequences, leveraging both active power and reactive power features.
Seq2Point: A neural network architecture that maps a sequence of aggregate power readings to a point estimate of an individual appliance’s state (Zhang et al., 2018). Our implementation used a convolutional neural network with five convolutional layers:
@@
CNN-LSTM: A hybrid deep learning model combining Convolutional Neural Networks (CNNs) for feature extraction and Long Short-Term Memory (LSTM) networks for temporal sequence modeling (Kaselimi et al., 2020). The CNN component consisted of:
@@
Transformer: A deep learning model based on the self-attention mechanism, allowing for efficient sequence modeling without relying on recurrence (Vaswani et al., 2017). We implemented a Transformer encoder with the following parameters:
@@
UNetNILM: An adaptation of the U-Net architecture for NILM tasks, designed to capture multiscale temporal patterns in the data (Faustine et al., 2020). The model featured an encoder-decoder structure with skip connections:
@@
Seq2Seq: A sequence-to-sequence neural network model that learns to map input sequences of aggregate power to output sequences of disaggregated signals (Du et al., 2016). Our implementation used an encoder-decoder LSTM architecture with:
@@
DAE (Denoising Autoencoder): An autoencoder model trained to reconstruct the target signal from a corrupted version of the input, effectively isolating the micro-inverter’s contribution from the aggregate power (Jia et al., 2019). The model consisted of:
@@
BERT+ (Li et al., 2024) baseline following Li et al. (2024) architecture description on PV injection estimation task. We use 5-layer BERT blocks, a feed-forward dimension of 180, and positional encoding. Since BERT+ requires external weather data for PV estimation, we use: (i) Active power (P) from aggregated consumption with injection, (ii) Reactive power (Q), (iii) Time positional embeddings (hour_sin, hour_cos), and (iv) Weather features from NSRDB (GHI and Temperature).
SunDance (Chen and Irwin, 2017) for PV injection estimation; this method was originally designed for solar-only disaggregation from net meter data without appliance monitoring. It is a “black-box” method combining physical solar modeling with machine learning. We first construct a location-specific clear sky generation model using the geographical coordinates and NSRDB solar irradiance data. This establishes the theoretical maximum solar generation potential at each timestamp based on solar position and panel characteristics. Following the paper (Chen and Irwin, 2017), which identified a universal weather-solar effect, we train an SVM with an RBF kernel to map weather metrics to the fraction of maximum clear sky generation. The estimated PV is then subtracted from the aggregate signal. We use SVM-RBF kernel (C=100, gamma=’scale’) settings.
@@ Appendix E Full Experiment Results
E.1. Overview of All Results Cross All Datasets
@@
Table 6 and 7 show the overall average performance metrics on appliance state recognition and energy injection disaggregation tasks.
| Dataset | Method | Accuracy (%) | Recall (%) | Precision (%) | F1 (%) | |
| Laboratory Dataset (Real BTM Injection) | FHMM | 60.95 | 28.32 | 16.84 | 17.25 | |
| XGBoost | 85.82 | 48.20 | 52.10 | 47.75 | ||
| Seq2Point | 67.63 | 53.06 | 27.86 | 34.27 | ||
| CNN-LSTM | 74.03 | 52.97 | 30.29 | 36.67 | ||
| Transformer | 81.79 | 52.91 | 38.16 | 42.76 | ||
| UNetNILM | 89.20 | 54.01 | 60.81 | 53.89 | ||
| DualNILM | 98.54 | 85.43 | 84.16 | 84.54 | ||
| REDD (Synthetic PV Injection) | House 1 | FHMM | 79.52 | 1.46 | 13.47 | 2.07 |
| XGBoost | 97.90 | 74.50 | 88.74 | 79.18 | ||
| Seq2Point | 97.65 | 42.20 | 56.42 | 46.07 | ||
| CNN-LSTM | 98.67 | 46.32 | 50.75 | 47.96 | ||
| Transformer | 95.99 | 56.08 | 54.10 | 53.42 | ||
| UNetNILM | 99.05 | 76.11 | 85.22 | 75.27 | ||
| DualNILM | 99.84 | 96.26 | 97.11 | 96.64 | ||
| House 2 | FHMM | 75.08 | 0.58 | 0.34 | 0.43 | |
| XGBoost | 98.43 | 92.47 | 95.58 | 93.53 | ||
| Seq2Point | 91.43 | 52.90 | 58.44 | 50.85 | ||
| CNN-LSTM | 98.86 | 49.19 | 48.74 | 48.96 | ||
| Transformer | 97.92 | 47.78 | 47.98 | 47.87 | ||
| UNetNILM | 99.57 | 64.16 | 77.22 | 67.28 | ||
| DualNILM | 99.70 | 92.48 | 94.00 | 93.11 | ||
| House 3 | FHMM | 83.84 | 4.53 | 0.90 | 1.49 | |
| XGBoost | 94.89 | 46.22 | 56.38 | 50.51 | ||
| Seq2Point | 98.10 | 42.78 | 50.90 | 44.47 | ||
| CNN-LSTM | 95.92 | 28.85 | 38.17 | 32.62 | ||
| Transformer | 97.81 | 40.21 | 38.89 | 39.43 | ||
| UNetNILM | 99.07 | 59.91 | 54.73 | 57.10 | ||
| DualNILM | 99.69 | 77.42 | 84.46 | 80.04 | ||
| UKDALE (Synthetic PV Injection) | House 1 | FHMM | 89.25 | 1.58 | 0.76 | 1.01 |
| XGBoost | 93.64 | 41.74 | 68.04 | 45.97 | ||
| Seq2Point | 88.70 | 41.47 | 43.29 | 36.79 | ||
| CNN-LSTM | 98.33 | 31.76 | 31.29 | 31.52 | ||
| Transformer | 96.90 | 33.65 | 27.76 | 30.26 | ||
| UNetNILM | 98.76 | 52.03 | 72.01 | 55.20 | ||
| DualNILM | 99.72 | 95.23 | 94.30 | 94.71 | ||
| House 2 | FHMM | 82.95 | 0.62 | 0.52 | 0.55 | |
| XGBoost | 96.13 | 66.61 | 79.19 | 70.67 | ||
| Seq2Point | 98.83 | 42.68 | 46.49 | 41.92 | ||
| CNN-LSTM | 97.81 | 41.65 | 45.48 | 43.33 | ||
| Transformer | 99.13 | 59.88 | 77.44 | 61.73 | ||
| UNetNILM | 99.17 | 64.25 | 68.74 | 65.43 | ||
| DualNILM | 99.73 | 90.63 | 94.46 | 92.31 | ||
| Dataset | Method | Avg. RMSE | Avg. MAE | |
| Laboratory Dataset (Real Injection) | FHMM | 0.5015 | 0.4492 | |
| Seq2Seq | 0.3512 | 0.2622 | ||
| DAE | 0.8830 | 0.7851 | ||
| UNetNILM | 0.2305 | 0.1883 | ||
| DualNILM | 0.1429 | 0.0804 | ||
| REDD with Synthetic Injection | House 1 | FHMM | 0.8093 | 0.2200 |
| SunDance | 0.8523 | 0.2847 | ||
| Seq2Seq | 0.6834 | 0.2370 | ||
| DAE | 0.8115 | 0.2171 | ||
| BERT+ | 0.7234 | 0.1945 | ||
| UNetNILM | 0.7589 | 0.1983 | ||
| DualNILM | 0.6661 | 0.1777 | ||
| House 2 | FHMM | 0.8403 | 0.3367 | |
| SunDance | 0.9271 | 0.3542 | ||
| Seq2Seq | 0.7217 | 0.3287 | ||
| DAE | 0.8448 | 0.3356 | ||
| BERT+ | 0.7812 | 0.3012 | ||
| UNetNILM | 2.2077 | 0.6601 | ||
| DualNILM | 0.7065 | 0.2768 | ||
| House 3 | FHMM | 0.5575 | 0.1766 | |
| SunDance | 0.5678 | 0.2134 | ||
| Seq2Seq | 0.4974 | 0.1940 | ||
| DAE | 0.5606 | 0.1743 | ||
| BERT+ | 0.4723 | 0.1623 | ||
| UNetNILM | 2.2517 | 0.4708 | ||
| DualNILM | 0.4261 | 0.1422 | ||
| Avg. | FHMM | 0.7357 | 0.2444 | |
| SunDance | 0.7824 | 0.2841 | ||
| Seq2Seq | 0.6342 | 0.2532 | ||
| DAE | 0.7390 | 0.2423 | ||
| BERT+ | 0.6590 | 0.2193 | ||
| UNetNILM | 1.7394 | 0.4431 | ||
| DualNILM | 0.5996 | 0.1989 | ||
| UKDALE with Synthetic Injection | House 1 | FHMM | 1.1905 | 0.3331 |
| SunDance | 1.1523 | 0.3478 | ||
| Seq2Seq | 0.9055 | 0.2758 | ||
| DAE | 0.9197 | 0.2948 | ||
| BERT+ | 0.9712 | 0.2945 | ||
| UNetNILM | 7.5014 | 1.3871 | ||
| DualNILM | 0.9008 | 0.2731 | ||
| House 2 | FHMM | 1.1477 | 0.3521 | |
| SunDance | 1.2187 | 0.3687 | ||
| Seq2Seq | 0.9252 | 0.2833 | ||
| DAE | 0.9594 | 0.3368 | ||
| BERT+ | 1.0123 | 0.3123 | ||
| UNetNILM | 1.1285 | 0.3043 | ||
| DualNILM | 0.9344 | 0.2883 | ||
| Avg. | FHMM | 1.1691 | 0.3426 | |
| SunDance | 1.1855 | 0.3583 | ||
| Seq2Seq | 0.9153 | 0.2795 | ||
| DAE | 0.9395 | 0.3158 | ||
| BERT+ | 0.9917 | 0.3034 | ||
| UNetNILM | 4.3150 | 0.8457 | ||
| DualNILM | 0.9176 | 0.2807 | ||
E.2. Laboratory Dataset Results
Table 8 shows the details of the laboratory dataset performance metrics for DualNILM and benchmarks.